OpenStudy (anonymous):

WILL GIVE MEDALS!!!! Xavier sells homemade necklaces for a profit of $22 per necklace. He is considering switching to a new type of pendant that will increase his profit, as expressed by the function y = 29x, where x is the number sold and y is the amount of profit. How many more dollars will Xavier earn on each necklace if he switches to the new material? Explain how you found your answer.

OpenStudy (anonymous):

do you have any idea what you would do here? I am going to guess in (y=29x) 29 would be the dollar amount. 29*x 22*x

OpenStudy (anonymous):

Lets give x a number to make this equation, would you like to provide me with a number so I can tell if you are paying attention?

OpenStudy (anonymous):

Thanks or trying to help but I already figured it out...I think. I got that he would save 7 dollars. Is that right?

OpenStudy (anonymous):

@kathlyn98 Am I right? I used the number one

OpenStudy (anonymous):

I am not sure where my response is, so I will write again, lol 27-22=7 yes, but what is your response to how you figured it out?

OpenStudy (anonymous):

Yes, I said that I plugged 1 in the formula. Like this y=29*1=29. I already submitted it just waiting for my teacher to grade it. Thanks for your help.